SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Forward Endcap monthly plateau

  • We are routinely checking the status of the RPC installed in the Forward Endcap in 2002 by taking cosmics data once a month
    • It is one of the task of the IFR OpMan to schedule it

  • From Epics:
    1. Make sure we are in DENY mode for injection (check the BaBar State Machine)
    2. From the State Machine Panel place the detector into manual mode by pressing the change button under the HV MODE
    3. Open the IFR EPICS folder (making sure there are no alarms)
    4. Open the IFR configuration panel from the IFR EPICS folder
    5. Put the IFR in LOCAL mode by clicking on GLOBAL
    6. Set the IFR state to RUNNABLE

  • From ORC:
    1. Shutdown ORC and go to the STANDBY STATE
    2. From the ORC GUI, click the main control button and select RUN TYPE.  Set the run type to COSMICS_NODCH
    3. From the BaBar proxies main panel include the the following subdetectors: IFR + GLT + CEN:BIP
    4. Ignore that IFR is not runnable
    5. Click on Change trigger mask from the setting menu. The trigger mask is 1000221

  • From Epics:
    1. In the IFR configuration panel set the Run Volt cell for Forw. EC A (first column, row 0) to 6100V and the Run Volt cell for Forw Aval. A (first column, row 8) to 8700V. (we are changing voltage for EVEN channels)
    2. Loop on the following operations:
      1. Wait until the voltages are ramped to the new settings (look into HV Crate #4 for Streamer and HV Crate #5 for Avalanche mode RPC)
      2. From ORC: Begin a new run
      3. Rate depends on the voltage and should be around 2.5-5 kHz. Take ~200.000 events.
      4. Take a note of the Voltage, Run Number, Rate, Number of events
      5. From ORC: Stop the run
      6. From Epics: in the IFR configuration panel increase the voltage of the channels under study by 200 Volts.
        The last point is 7100V for Streamer RPC and 9600V for Avalanche RPC (6 runs in total)
    3. When last point has been taken, set the voltage of the studied channels to RUNNBLE (i.e.: 6700V for streamer and 9200V for Avalanche)

  • Loop again on the same steps but with ODD channels
    (i.e.: change the voltage in steps of 200V starting from (6100V, 8600V) up to (7100V, 9600V) for (Forw. EC B, Forw. Aval. B)


  • When finished, from ORC:
    1. Return to the STANDBY STATE and then restore the default trigger mask by clicking on CONTROL+ and then RESTORE_DEFAULTS. Change the run type back to PHYSICS before starting a new run.

  • From Epics:
    1. Make sure that you have set the Run Voltage to (6700V,9200V) for (Forw EC. A(B), Forw. Aval. A(B))
    2. Set IFR to INJECTABLE
    3. Return IFR into GLOBAL mode by clicking on LOCAL
    4. Double check that all the detectors are rumped down before permetting the injection
    5. Insert a note of your activities in the logbook and post a hypernews on IFR Operations

  • For any problems on the calibration setup contact the IFR DAQ experts.

  • Useful stuff:
  • EVEN/ODD channels described here
  • Old instructions for streamer and for avalanche


  • Last modified: 2 Sept 2007
    by Antonio Petrella
    For any question, please contact: